# Contextual Gesture: Co-Speech Gesture Video Generation through Context-aware Gesture Representation

### v1

@@ 3.3 Structure-Aware Image Generation
Edge Heatmaps.
Using the gesture motion keypoints, we establish linkages between them to provide structural information. To optimize computational efficiency, we limit the number of keypoint connections to those defined by body joint relationships (Wan et al., 2017), rather than all potential connections in (He et al., 2023).
@@
Structural-guided Image Refinement.
Traditional optical-flow-based warping methods are effective for handling global deformations but often fail under large motion patterns, such as those involving hands or shoulders, resulting in significant distortions. To address this, we introduce a structure-guided refinement process that incorporates semantic guidance via structural heatmaps.
Instead of directly rendering the warped feature maps into RGB images, we first predict a low-resolution feature map of size . Multi-resolution edge heatmaps are generated and used as structural cues to refine the feature maps. After performing deformation and occlusion prediction at each scale using TPS (Zhao & Zhang, 2022), the edge heatmaps are fed into the generator. Specifically, we integrate these heatmaps into the fusion block using SPADE (Park et al., 2019), enabling the precise prediction of residuals. These residuals are element-wise added to the warped feature maps, ensuring precise structural alignment.
@@
We employ an adversarial loss, along with perceptual similarity loss (LPIPS) (Johnson et al., 2016) and pixel-level loss for image refinement. The reconstruction objective is defined as:
@@
where denotes the ground-truth image, and represents the generated image. We leverage a small weighted term of to prevent GAN training collapse.
4 Experiments
Since our work focuses on joint gesture motion and video generation, the main experiments primarily compare our approach with existing methods that also address joint generation. Comparisons for gesture motion generation and avatar video rendering are deferred to the Appendix.
4.1 Experimental Settings
@@
Dataset and Preprocessing. We utilize PATS (Ginosar et al., 2019; Ahuja et al., 2020) for the experiments. It contains 84,000 clips from 25 speakers with a mean length of 10.7s, 251 hours in total. For a fair comparison, following the literature (Liu et al., 2022c; He et al., 2024) and replace the missing subject, with 4 speakers are selected (Noah, Kubinec, Oliver, and Seth). All video clips are cropped with square bounding boxes, centering speaks, resized to . We defer the additional details in the Appendix.
4.2 Gesture Video Generation
Evaluation Metrics.
@@ 4.2 Gesture Video Generation
For gesture motion metrics, we use Fréchet Gesture Distance (FGD) (Yoon et al., 2020) to measure the distribution gap between real and generated gestures in feature space, Diversity (Div.) (Lee et al., 2019) to calculate the average feature distance between generated gestures, Beat Alignment Score (BAS) following (Li et al., 2021), and Percent of Correct Motion parameters (PCM), difference of generation deviate from ground-truth following (Chen et al., 2024). We extract 2D human poses for face and body using MMPose (OpenMMLab, 2020). Note that in comparison with other models, the FGD is measued by the keypoints extracted from generated videos while in ablation studies, to prevent the effect image warping errors, FGD is measured by keypoints generated in Sec.3.2.
@@
For pixel-level video quality, we assess Fréchet Video Distance (FVD) (Unterthiner et al., 2018) for the overall quality of gesture videos, VQAA for aesthetics and VQAT for technical quality based on Dover (Wu et al., 2023), pretrained on datasets with labels ranked by real users.
We further evaluate the training and inference efficiency of various methods, Train-T denotes the number of days for training. Infer-T denotes the number of seconds to produce a 10-second video.
@@
| MM-Diffusion | 67.56 | 4.32 | 0.65 | 0.11 | - | 77.65 | 4.14 | 14 days | 600 sec |
| ANGIE | 34.13 | 7.87 | 0.78 | 0.37 | 515.43 | 86.32 | 4.98 | 5 days | 30 sec |
| S2G-Diffusion | 10.54 | 10.08 | 0.98 | 0.45 | 493.43 | 94.54 | 5.63 | 5 days | 35 sec |
| EchoMimicV2 | 13.65 | 9.85 | 0.98 | 0.45 | 466.84 | 95.65 | 5.98 | - | 1200 sec |
| Ours + AnimateAnyone | 6.56 | 13.06 | 0.99 | 0.54 | 477.82 | 95.63 | 6.04 | 8.5 days | 50 sec |
@@
Baseline Methods.
We benchmark Contextual Gesture against several co-speech gesture video generation methods: (1) ANGIE (Liu et al., 2022c), (2) S2G-Diffusion (He et al., 2024), (3) MM-Diffusion (Ruan et al., 2023), and (4) EchoMimicV2 (Meng et al., 2024). The first two are conventional optical-flow-based methods. MM-Diffusion is capable of achieving joint audio-visual generation. EchoMimicV2 is the most recent diffusion based speech-avatar animation model pretrained on large scale data.
Evaluation Results.
We present quantitative evaluations in Tab. 1. Our approach significantly outperforms existing methods in both gesture motion and video quality metrics. We provide qualitative evaluations in Fig. 4. MM-Diffusion is not able to handle complex motion patterns, leading to almost static results. ANGIE and S2G-Diffusion struggle with local regions, such as the hands, due to its reliance on unsupervised keypoints for global transformations, which neglects local deformations. EchoMimicV2 lacks the background motion modeling and is only capable of presenting the aligned centered avatars in the middle. In addition, it fails to achieve diversified gestures. In contrast, our method demonstrates high-quality video generation, particularly in the facial and body areas. The alignment between gesture and speech is notably enhanced through our speech-content-aware gesture latent representation. For example, when the actor says ”90 joules,” he points to the screen, and he emphasizes phrases like ”so two ways” and ”in each case” by raising his hands.
User Study.
We conducted a user study to evaluate the visual quality of our method. We sampled 80 videos from each method including EchoMimicV2, S2G-Diffusion, ANGIE and ours and invited 20 participants to conduct Mean Opinion Scores (MOS) evaluations. The rating ranges from 1 (poorest) to 5 (highest). Participants rated the videos on: (1) MOS1: “How realistic does the video appear?”, (2) MOS2: “How diverse does the gesture pattern present?”, (3) MOS3: “Are speech and gesture synchronized in this video?”. The videos were presented in random order to capture participants’ initial impressions. As shown in Fig. 5, our method outperformed others across realness, synchronization and diversity, achieving significant performance improvement over existing methods.
@@
4.3 Ablation Study
Motion Keypoint Design. We evaluate four settings for image-warping: (1) unsupervised keypoints for global optical-flow transformation (as in ANGIE and S2G-Diffusion), (2) 2D human poses, (3) 2D human poses augmented with flexible learnable points, and (4) full-model reconstruction with refinement. Each design is assessed using TPS (Zhao & Zhang, 2022) transformation, with self-reconstruction based on these keypoints for evaluation. As shown in Tab. 2(a), learnable keypoints lead to a significant decrease in FVD, highlighting their inadequacy for motion control. The inclusion of flexible keypoints does not enhance the image-warping outcomes. Consequently, we opt to utilize 2D pose landmarks exclusively for our study.
4.4 Application
Long Sequence Generation.
As in Fig. 7, to generate long sequences, we begin with the initial frame and corresponding target audio, segmented into smaller windows. After generating the first segment, the last few frames of the output serve as the new starting condition for the next segment, enabling iterative outpainting.
Video Gesture Editing.
For editing, we extract keypoints from the video, tokenize face and body movements into motion tokens, and insert mask tokens where edits are needed. By changing the speech audio or speaker identity, we can create new gesture patterns and re-render the video.
Gesture Pattern Transfer.
With different identity embeddings, we generate unique gesture patterns for the same audio input. See the demo videos in the Appendix.
Speech-Gesture Retrieval.
With chronological speech-gesture alignment, the model is capable of retrieving the best gesture motion correponding to the given speech audio in a batch of data. See additional details in the Appendix.
5 Conclusion
We present Contextual Gesture, a framework for generating realistic co-speech gesture videos. To ensure the gestures cohere well with speech, we propose speech-content aware gesture motion representation though knowledge distillation from the gesture-speech aligned features. Our structural-aware image generation module improves the transformation of latent motions into realistic animations for large-scale body motions. We hope this work encourage further exploration of the relationship between gesture patterns and speech context for more compelling gesture video generations in the future.
6 Impact Statements
This work focuses on generating co-speech gesture videos, highlighting both its potential and ethical concerns, particularly the risk of misuse for creating deepfakes or fabricated events involving public figures. While it’s impossible to fully eliminate such risks, the research provides a technical analysis of gesture video synthesis, emphasizing its capabilities and limitations. To mitigate misuse, we recommend practices like watermarking videos and using detection tools for synthetic content to promote responsible and ethical use of the technology.
@@ Appendix B Additional Related Works
Masked Representation Learning has been demonstated an effective representation learning for various modalities. (Devlin, 2018; He et al., 2022) Some works explored the generation capabilities using this paradigm. MAGE (Li et al., 2023) achieves high-quality image generation through iterative remasking. Muse (Chang et al., 2023) extends this idea to leverage language with region masking for image editing and achieve fine-grained control. Recent Masking Models (Pinyoanuntapong et al., 2024; Guo et al., 2024a; Wang, 2023; Mao et al., 2024) bring this strategy to the motion and gesture domain and improves the motion generation speed, quality, and editing capability. Inspired by these work, we propose the masked gesture generation conditioned the audio to learn the gesture-speech correspondence during generation.
@@ Appendix C Dataset Details and Preprocessing
C.1 Preprocessing
We found that many videos used in ANGIE (Liu et al., 2022c) and S2G-Diffusion (He et al., 2024), particularly for the subject Jon, are no longer available. To address this, we replaced Jon with Noah. We utilized the PATS (Ginosar et al., 2019) metadata to download videos from YouTube and preprocess them. After filtering, we obtained 1080 videos for Oliver, 1080 for Kubinec, 1080 for Seth, and 988 for Noah. For the testing dataset, we collected 120 videos for Oliver, 120 for Kubinec, 120 for Seth, and 94 for Noah.
@@ C.2 Feature Representation
C.3 Dataset License.
The video data within PATS dataset include personal identity information, and we strictly adhere to the data usage license “CC BY - NC - ND 4.0 International,” which permits non-commercial use.
@@ Appendix D Additional Implementation Details
We jointly train the framework on three speakers. The following sections provide the details for each module’s training.
@@
For pixel-level motion generation, we utilize Thin Plate Splines (TPS) (Zhao & Zhang, 2022). Our framework tracks 116 keypoints (68 for the face and 48 for the body). The number of TPS transformations is set to 29, with each transformation utilizing paired keypoints. In accordance with TPS methodologies, both the dense motion network and occlusion-aware generators leverage 2D convolutions to produce weight maps for optical flow generation, along with four occlusion masks at various resolutions (32, 64, 128, and 256) to facilitate image frame synthesis.

### v2

@@ 3.3. Structure-Aware Image Generation
Edge Heatmaps. Using the gesture motion keypoints, we establish linkages between them to provide structural information. To optimize computational efficiency, we limit the number of keypoint connections to those defined by body joint relationships (Wan et al., 2017), rather than all potential connections in (He et al., 2023).
@@
Structural-guided Image Refinement. Traditional optical-flow-based warping methods are effective for handling global deformations but often fail under large motion patterns, such as those involving hands or shoulders, resulting in significant distortions. To address this, we introduce a structure-guided refinement process that incorporates semantic guidance via structural heatmaps.
Instead of directly rendering the warped feature maps into RGB images, we first predict a low-resolution feature map of size . Multi-resolution edge heatmaps are generated and used as structural cues to refine the feature maps. After performing deformation and occlusion prediction at each scale using TPS (Zhao and Zhang, 2022), the edge heatmaps are fed into the generator. Specifically, we integrate these heatmaps into the fusion block using SPADE (Park et al., 2019) and the prediction of residuals are element-wise added to the warped feature maps, ensuring precise structural alignment.
@@
Training Objective. We employ an adversarial loss, along with perceptual similarity loss (LPIPS) (Johnson et al., 2016) and pixel-level loss for image refinement. The reconstruction objective is defined as:
@@
where and represent the ground-truth and generated image separately. We use a small weighted term of to stablize training.
4. Experiments
Since our work focuses on joint gesture motion and video generation, to validate the design, we first compare our proposed method for gesture motion generation with relevant 3D gesture motion generation frameworks. We further conduct holistic co-speech gesture video generation comparisons.
4.1. 3D Gesture Motion Generation.
Dataset. We select BEAT-X (Liu et al., 2023) as the dataset for comparison of gesture generation. For consistency, we exclude the image-to-animation component from our method and leverage 3D SMPL-X poses as in the existing literature. We compare the gesture generation module of our work with representative state-of-the-art methods in co-speech gesture generation (Ao et al., 2022; Yi et al., 2023; Liu et al., 2023). We further design a baseline without using contextual distillation.
Evaluation Metrics We evaluate the realism of body gestures by Fréchet Gesture Distance (FGD)(Yoon et al., 2020). We include Diversity by calculating the average L1 distance across clips. For synchronization, we use Beat Constancy (BC) (Li et al., 2021). For facial expression, we use the vertex Mean Squared Error (MSE) (Xing et al., 2023) for positional accuracy.
Experiment Results. As shown in table 1, our method significantly improves SMPL-X-based co-speech gesture generation, achieving lower FGD, higher diversity. These results indicate that our method produces smoother and more natural gesture motion patterns compared to existing approaches. Furthermore, the lower MSE combined with appropriate BC scores indicates that the generated gestures closely track the ground-truth gestures across frames while maintaining rhythmic alignment with the audio. This demonstrates that the gestures are well-coordinated with, and responsive to, the accompanying speech. These findings demonstrate the effectiveness of our contextual distillation strategy for motion representation learning, as well as the benefit of chronological alignment through contrastive learning. We defer the video comparisons in the Appendix videos for reference.
| Methods | FGD | BC | Div. | MSE |
| GT | 0.703 | 11.97 | ||
| HA2G (Liu et al., 2022c) | 12.320 | 0.677 | 8.626 | - |
| DisCo (Liu et al., 2022a) | 9.417 | 0.643 | 9.912 | - |
| CaMN (Liu et al., 2022d) | 6.644 | 0.676 | 10.86 | - |
| DiffSHEG (Chen et al., 2024) | 7.141 | 0.743 | 8.21 | 9.571 |
| TalkShow (Yi et al., 2023) | 6.209 | 0.695 | 13.47 | 7.791 |
| Rhythmic Gesticulator (Ao et al., 2022) | 6.453 | 0.665 | 9.132 | |
| EMAGE (Liu et al., 2023) | 5.512 | 0.772 | 13.06 | 7.680 |
| Ours (w/o distill) | 5.079 | 0.737 | 13.24 | 7.742 |
| Ours | 4.434 | 0.724 | 13.76 | 7.021 |
4.2. Realistic Video Generation.
@@ 4. Experiments
Dataset. We utilize PATS (Ginosar et al., 2019; Ahuja et al., 2020) for the experiments. It contains 84,000 clips from 25 speakers with a mean length of 10.7s, 251 hours in total. For a fair comparison, following the literature (Liu et al., 2022b; He et al., 2024) and replace the missing subject, with 4 speakers are selected (Noah, Kubinec, Oliver, and Seth). All video clips are cropped with square bounding boxes, centering speaks, resized to . We defer the additional details in the Appendix.
@@
Evaluation Metric. For gesture motion metrics, we use Fréchet Gesture Distance (FGD) (Yoon et al., 2020) to measure the distribution gap between real and generated gestures in feature space, Diversity (Div.) (Lee et al., 2019) to calculate the average feature distance between generated gestures, Beat Alignment Score (BAS) following (Li et al., 2021), Percent of Correct Motion parameters (PCM), difference of generation deviate from ground-truth following (Chen et al., 2024). For video generation, we extract 2D human poses for face and body using MMPose (OpenMMLab, 2020) to represent the gesture motion. Note that in comparison with other models, the FGD is measued by the keypoints extracted from generated videos in the main experiment while in ablation studies, to prevent the effect image warping errors, FGD is measured by keypoints generated in Sec.3.2.
@@
For pixel-level video quality, we assess Fréchet Video Distance (FVD) (Unterthiner et al., 2018) for the overall quality of gesture videos, VQAA for aesthetics and VQAT for technical quality based on Dover (Wu et al., 2023), pretrained on datasets with labels ranked by real users. We further evaluate the training and inference efficiency of various methods, Train-T denotes the number of days for training. Infer-T denotes the number of seconds to produce a 10-second video.
Baseline Methods. We benchmark Contextual Gesture against several co-speech gesture video generation methods: (1) ANGIE (Liu et al., 2022b), (2) S2G-Diffusion (He et al., 2024), (3) MM-Diffusion (Ruan et al., 2023), and (4) EchoMimicV2 (Meng et al., 2024). The first two are conventional optical-flow-based methods. MM-Diffusion is capable of achieving joint audio-visual generation. EchoMimicV2 is the most recent diffusion based speech-avatar animation model pretrained on large scale data.
Evaluation Results. We present quantitative evaluations in table 2. Our approach significantly outperforms existing methods in both gesture motion and video quality metrics. We provide qualitative evaluations in fig. 4. MM-Diffusion is not able to handle complex motion patterns, leading to almost static results. ANGIE and S2G-Diffusion struggle with local regions, such as the hands, due to its reliance on unsupervised keypoints for global transformations, which neglects local deformations. EchoMimicV2 lacks the background motion modeling and is only capable of presenting the aligned centered avatars in the middle. In addition, it fails to achieve diversified gestures. In contrast, our method demonstrates high-quality video generation, particularly in the facial and body areas. The alignment between gesture and speech is notably enhanced through our speech-content-aware gesture latent representation. For example, when the actor says ”90 joules,” he points to the screen, and he emphasizes phrases like ”so two ways” and ”in each case” by raising his hands.
@@
| MM-Diffusion (Ruan et al., 2023) | 67.56 | 4.32 | 0.65 | 0.11 | - | 77.65 | 4.14 | 14 days | 600 sec |
| ANGIE (Liu et al., 2022b) | 34.13 | 7.87 | 0.78 | 0.37 | 515.43 | 86.32 | 4.98 | 5 days | 30 sec |
| S2G-Diffusion (He et al., 2024) | 10.54 | 10.08 | 0.98 | 0.45 | 493.43 | 94.54 | 5.63 | 5 days | 35 sec |
| EchoMimicV2 (Meng et al., 2024) | 13.65 | 9.85 | 0.98 | 0.45 | 466.84 | 95.65 | 5.98 | - | 1200 sec |
| Ours + AnimateAnyone(Hu et al., 2023) | 6.56 | 13.06 | 0.99 | 0.54 | 477.82 | 95.63 | 6.04 | 8.5 days | 50 sec |
@@
User Study. We conducted a user study to evaluate the visual quality of our method. We sampled 80 videos from each method including EchoMimicV2, S2G-Diffusion, ANGIE and ours and invited 20 participants to conduct Mean Opinion Scores (MOS) evaluations. The rating ranges from 1 (poorest) to 5 (highest). Participants rated the videos on: (1) MOS1: “How realistic does the video appear?”, (2) MOS2: “How diverse does the gesture pattern present?”, (3) MOS3: “Are speech and gesture synchronized in this video?”. The videos were presented in random order to capture participants’ initial impressions. As shown in fig. 5, our method outperformed others across realness, synchronization and diversity, achieving significant performance improvement over existing methods.
4.3. Ablation Study
Motion Keypoint Design. We evaluate four settings for image-warping: (1) unsupervised keypoints for global optical-flow transformation (as in ANGIE and S2G-Diffusion), (2) 2D human poses, (3) 2D human poses augmented with flexible learnable points, and (4) full-model reconstruction with refinement. Each design is assessed using TPS (Zhao and Zhang, 2022) transformation, with self-reconstruction based on these keypoints for evaluation. As shown in table 3(a), learnable keypoints lead to a significant decrease in FVD, highlighting their inadequacy for motion control. The inclusion of flexible keypoints does not enhance the image-warping outcomes. Consequently, we opt to utilize 2D pose landmarks exclusively for our study.
Training Strategies. We evaluate the mask ratio during training and the number of inference steps during decoding. As shown in table 3(d), our model requires only 5 inference steps, in contrast to over 50 or 100 steps in diffusion-based models. Furthermore, a uniform masking ratio between 0.5 and 1 during training yields optimal performance.
Long Sequence generation. To understand the capability of our framework for long sequence generation, we conduct an ablation study for both PATS and BEAT-X dataset. For BEAT-X, we cut the testing audios into segments of 256 (about 8.53 seconds) for short sequence evaluation and use raw testing audios for long sequence evaluation in table 1. Shown in table 3(f), it is interesting for PATS dataset, long-sequence generation as an application in the main paper presents quality lower than normal settings. However, for BEAT-X dataset, the generation quality is not affected much. We attribute this difference caused by the dataset difference. Because PATS dataset consists training video lengths with a average of less than 10 seconds, the model presents less diverse gesture patterns. However, in BEAT-X, most of gesture video sequences are over 30 or 1 minutes, our method further benefits from this long sequence learning precess and presents higher qualities.
@@ 4.3. Ablation Study
| iter. | FGD | Div. | PCM |
| 5 | 0.87 | 13.23 | 0.59 |
| 10 | 0.98 | 13.11 | 0.57 |
| 15 | 1.24 | 13.04 | 0.57 |
| 20 | 1.56 | 13.23 | 0.57 |
| M-Ratio | FGD | Div. | PCM |
| Uni 0-1 | 2.13 | 14.31 | 0.56 |
| Uni .3-1 | 1.56 | 12.44 | 0.512 |
| Uni .5-1 | 0.87 | 13.23 | 0.59 |
| Uni .7-1 | 1.22 | 13.12 | 0.57 |
| Dataset | Setting | FGD | Div. | BAS |
| PATS | 10s | 1.303 | 13.260 | 0.996 |
| 10s | 2.356 | 11.956 | 0.994 | |
| BEAT-X | 10s | 4.747 | 13.14 | 7.323 |
| 10s | 4.650 | 13.55 | 7.370 |
@@
4.4. Application
Long Sequence Generation. As in fig. 7, to generate long sequences, we begin with the initial frame and corresponding target audio, segmented into smaller windows. After generating the first segment, the last few frames of the output serve as the new starting condition for the next segment, enabling iterative outpainting.
Video Gesture Editing. For editing, we extract keypoints from the video, tokenize face and body movements into motion tokens, and insert mask tokens where edits are needed. By changing the speech audio or speaker identity, we can create new gesture patterns and re-render the video.
Gesture Pattern Transfer. With different identity embeddings, we generate unique gesture patterns for the same audio input. See the demo videos in the Appendix.
Speech-Gesture Retrieval. With chronological speech-gesture alignment, the model is capable of retrieving the best gesture motion correponding to the given speech audio in a batch of data. See additional details in the Appendix.
5. Conclusion
We present Contextual Gesture, a framework for generating realistic co-speech gesture videos. To ensure the gestures cohere well with speech, we propose speech-content aware gesture motion representation though knowledge distillation from the gesture-speech aligned features. Our structural-aware image generation module improves the transformation of latent motions into realistic animations for large-scale body motions. We hope this work encourage further exploration of the relationship between gesture patterns and speech context for better video generations in the future.
@@ Appendix B Additional Related Works
Masked Representation Learning has been demonstated an effective representation learning for various modalities. (Devlin, 2018; He et al., 2022) Some works explored the generation capabilities using this paradigm. MAGE (Li et al., 2023) achieves high-quality image generation through iterative remasking. Muse (Chang et al., 2023) extends this idea to leverage language with region masking for image editing and achieve fine-grained control. Recent Masking Models (Pinyoanuntapong et al., 2024; Guo et al., 2024a; Wang, 2023; Mao et al., 2024; Zhang et al., 2025) bring this strategy to the motion and gesture domain and improves the motion generation speed, quality, and editing capability. Inspired by these work, we propose the masked gesture generation conditioned the audio to learn the gesture-speech correspondence during generation.
@@ Appendix C Dataset Details and Preprocessing
C.1. Preprocessing
We found that many videos used in ANGIE (Liu et al., 2022b) and S2G-Diffusion (He et al., 2024), particularly for the subject Jon, are no longer available. To address this, we replaced Jon with Noah. We utilized the PATS (Ginosar et al., 2019) metadata to download videos from YouTube and preprocess them. After filtering, we obtained 1080 videos for Oliver, 1080 for Kubinec, 1080 for Seth, and 988 for Noah. For the testing dataset, we collected 120 videos for Oliver, 120 for Kubinec, 120 for Seth, and 94 for Noah.
@@ C.2. Feature Representation
C.3. Dataset License.
The video data within PATS dataset include personal identity information, and we strictly adhere to the data usage license ``CC BY - NC - ND 4.0 International,'' which permits non-commercial use.
@@ Appendix D Additional Implementation Details
We jointly train the framework on three speakers. The following sections provide the details for each module's training.
@@
For pixel-level motion generation, we utilize Thin Plate Splines (TPS) (Zhao and Zhang, 2022). Our framework tracks 116 keypoints (68 for the face and 48 for the body). The number of TPS transformations is set to 29, with each transformation utilizing paired keypoints. In accordance with TPS methodologies, both the dense motion network and occlusion-aware generators leverage 2D convolutions to produce weight maps for optical flow generation, along with four occlusion masks at various resolutions (32, 64, 128, and 256) to facilitate image frame synthesis.